OpenStudy (anonymous):

jeff is playing cards with a friend. his opponent just dealt him an almost perfect hand all he needs to win is a five. they each have 8 cards. if his opponent doesnt have a five, and jeff has 2 fives, what is the probability that the card jeff draws is a five? ?/?

OpenStudy (anonymous):

how many cards are left? how many of them are a 5?

OpenStudy (anonymous):

Each of them have 8 cards, so there are 52-3*8 = 28 cards left in the deck. Jeff has two of the fives already, and his opponent has none, so there are still 2 fives left. If he draws a card, there's a 2/28 = 1/14 chance he draws the card he needs.

OpenStudy (anonymous):

there's only 2 people playing, i believe

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Sorry, don't know how I got 3 into my head. You can redo the calculations easily for 2 players.

OpenStudy (anonymous):

ain't no thing but a chicken wing